# Balancing Act: Distribution-Guided Debiasing in Diffusion Models

### v1

@@ Abstract
Diffusion Models (DMs) have emerged as powerful generative models with unprecedented image generation capability. These models are widely used for data augmentation and creative applications. However, DMs reflect the biases present in the training datasets. This is especially concerning in the context of faces, where the DM prefers one demographic subgroup vs others (eg. female vs male). In this work, we present a method for debiasing DMs without relying on additional data or model retraining. Specifically, we propose Distribution Guidance, which enforces the generated images to follow the prescribed attribute distribution. To realize this, we build on the key insight that the latent features of denoising UNet hold rich demographic semantics, and the same can be leveraged to guide debiased generation. We train Attribute Distribution Predictor (ADP) - a small mlp that maps the latent features to the distribution of attributes. ADP is trained with pseudo labels generated from existing attribute classifiers. The proposed Distribution Guidance with ADP enables us to do fair generation. Our method reduces bias across single/multiple attributes and outperforms the baseline by a significant margin for unconditional and text-conditional diffusion models. Further, we present a downstream task of training a fair attribute classifier by rebalancing the training set with our generated data. Project Page
@@ 1 Introduction
Recent advancements in Diffusion Models (DM) [11, 33, 42] have garnered much interest among researchers in evaluating the quality of the generated content. These models are not only used to generate realistic content but also to augment real datasets [45, 24] for downstream tasks. However, these popular models have been found to synthesize biased content with respect to multiple demographic factors like gender, race, age, etc., which can be detrimental to society once these models are deployed in the real world [29, 21, 34]. The problem is largely caused by the images used to train these models, as the outputs of these models are governed by these training datasets [29, 22]. Effects of such harmful biases have been shown by multiple recent works involving studies on not only DMs [29, 21, 34], but also GANs and other generative models [22, 5, 12]. In fact, Perera et al. [29] show that unconditional DMs–even when trained with balanced data–amplify racial biases, leading to the generation of more white-skinned faces than dark-skinned ones. Biases in the generated data are even more evident in large text-to-image DMs, e.g., models mostly tend to generate a specific gender with a given profession (like male and doctor) [23, 52, 39].
@@
Existing works on debiasing, either require a reference dataset  [5, 52] and/or allow retraining of the model [5, 50, 51]. On the contrary, our work aims to mitigate biases in both unconditional and conditional DMs, enabling fair generation without model retraining. We propose a setting where we provide a reference attribute distribution ($\mathbf{p^{a}_{ref}}$) for sensitive attribute $a$ and query the DM to sample images following $\mathbf{p^{a}_{ref}}$. This framework allows for adapting any existing DM to a pre-defined distribution. E.g., a user can provide $\mathbf{p^{a}_{ref}}$ as a uniform histogram of sensitive attributes to generate balanced attribute distribution. We believe defining $\mathbf{p^{a}_{ref}}$ provides just enough information to condition the DM for fair generation. This is an extremely practical setting for debiasing and is particularly important for large text-to-image DMs [33, 36, 32] where retraining or fine-tuning is computationally intensive.
@@
One plausible approach for fair generation is to guide every generated sample with attribute classifiers following classifier guidance [8]. However, such a framework, though simple, is overly restrictive as it requires presetting and enforcing attributes for each sample individually (which we call sample guidance). Such constraints during denoising result in inferior generation quality, as discussed in Sec.4.1. Instead, we propose to jointly denoise a whole batch of samples and guide the process with $\mathbf{p^{a}_{ref}}$ (which we call distribution guidance). Specifically, we push the generated batch attribute distribution $\mathbf{{p^{a}_{\theta}}}$ and $\mathbf{p^{a}_{ref}}$ close to each other during the reverse process. Distribution guidance provides more flexibility to each sample during generation as it does not enforce a preset of attributes on a sample basis. This results in fair generation without sacrificing the generation quality.
@@
A major challenge for guidance-based conditioning is that it requires separate image classifiers for each noise scale of the diffusion process. To overcome this, we propose to perform guidance in semantically rich features space - h-space [18] of DMs. Specifically, we train an Attribute Distribution Predictor (ADP) that predicts attribute distribution directly from the h-space features. As ADP is trained on rich and discriminative h-space features, it is - a) implemented as a linear layer, b) requires minimal training data, and c) fast training and inference. Finally, during inference, we steer the h-space representation by matching the predictions from ADP to $\mathbf{p^{a}_{ref}}$.
@@
    A novel setting for debiasing existing DMs without retraining given a reference attribute distribution.
@@ 2 Related Works
Biases in Generative Models. While generative models like Generative Adversarial Networks and Diffusion Models have become the de-facto tools for image generation in recent times, studies show that they are not free of biases [29, 21, 34, 22, 13]. Perera et al.  [29] show that unconditional diffusion models amplify the biases in the training data with respect to gender, race, and age. Luccioni et al.  [21] identify and quantify social biases in images generated by popular text-to-image models like DALL-E 2, and Stable Diffusion v1.4 and 2. Maluleke et al.  [22] study racial biases in GANs and find that GANs mimic the racial distribution in the training data.
@@
Debiasing generative models without training. As training of GANs and DMs can be resource-consuming, many methods prefer fair generation of images without explicit training [31, 43, 24, 12]. MaGNET [12] aims to produce uniform sampling on the learned manifold of any generative model like GANs or VAEs, while Ramaswamy et al.  [31] and Tan et al.  [43] manipulate the latent space of GANs to generate balanced outputs. GANDiffFace [24], on the other hand, generates balanced synthetic for face recognition by first generating high-quality images from different demographics using GANs and then finetuning Stable Diffusion [33] using DreamBooth [35] to generate more images of such identities with different poses, expressions, etc. Multiple works attempt to mitigate biases in vision-language models and text-conditioned diffusion models as well [6, 52, 41, 53, 3]. Chuang et al.  [6] debias the text embedding using a calibrated projection matrix and shows unbiased generations without any additional training or data. However, debiasing the unconditional DMs has received less attention, which is the main focus of this work.
@@ 3 Method
We assume a setting where we are given a pretrained DM trained on biased data and a reference distribution of the sensitive attributes $\mathbf{p^{a}_{ref}}$. Our goal is to generate data from the DM, whose generated attribute distribution $\mathbf{p^{a}_{\theta}}$ best approximates the reference $\mathbf{p^{a}_{ref}}$ without retraining.
The key idea is to jointly guide the denoising of a batch of samples such that $\mathbf{p^{a}_{ref}}$approx $\mathbf{p^{a}_{\theta}}$. Directly computing $\mathbf{p^{a}_{\theta}}$ in a closed form is intractable. Instead, we train an Attribute Distribution Predictor, a small network that maps the intermediate batch features of a diffusion network to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
Figure 2: Sample guidance vs Distribution guidance. a) After a few steps of denoising ($t=\tau$), the generated samples have learned some discriminative features for gender. Sample guidance enforces a preset attribute state (male/female) to all the samples in the batch. This is stringent and can result in enforcing samples with dominant female features to be changed. However, in distribution guidance, only those samples are converted, which are easy to convert and are close to the decision boundary. b) Visualization of the generated samples with different guidance strengths. It can be observed that sample guidance is not able to change the attribute in some cases, and increasing guidance strength results in collapse.
@@ 3.1 Preliminary
Diffusion models have emerged as a powerful family of generative models trained to learn the data distribution by gradual denoising from a Gaussian distribution. Starting from a clean point $\mathbf{x_{0}}$, and a set of scalar values $\{\alpha_{t}\}_{t=1}^{T}$, applying $\mathbf{t}$ steps of the forward diffusion process yields a noisy data point, where $\bar{\alpha_{t}}$ = $\prod_{i=1}^{t}\alpha_{i}$
$\mathbf{x_{t}}=\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}}+(\sqrt{1-\bar{\alpha_{t}}})\epsilon,\epsilon\approx\mathcal{N}(0,I)$ (1)
A diffusion model is learned as a neural network $\epsilon_{\theta}$ that predicts the noise from given $\mathbf{x_{t}}$ and $\mathbf{t}$. The reverse process takes the form $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{x_{0}})$, which is parameterized as a Gaussian distribution. In this work, we consider DDIM [42] sampling which first computes a clean data point $\mathbf{\hat{x}_{0}}$ and then sample $\mathbf{x_{t-1}}$ from $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{\hat{x}_{0}})$.
Classifier guidance is proposed to condition a diffusion model on class labels with the help of a pretrained classifier [8]. Specifically, a classifier $f_{\phi}(c|\mathbf{x_{t}},t)$ is trained on noisy images at different noise scales, where $c$ is the class label. The gradients of the classifier are then used to guide the diffusion sampling process towards the class label $c$. Concretely, the classifier guidance performs sampling by updating the noise prediction $\epsilon_{\theta}(\mathbf{x_{t}},\mathbf{t})$.
@@ 3.2 Classifier guidance for debiasing
A promising approach is to leverage pretrained attribute classifiers to guide toward balanced generation. Assuming a reference attribute distribution $\mathbf{p^{a}_{ref}}$ for a binary attribute $\mathbf{a}$ (e.g., gender) and corresponding attribute classifier ${f}_{\phi}$. We parameterize $\mathbf{p^{a}_{ref}}$ as a Bernoulli distribution with parameter $\mathbf{r}$ denoting fraction of males samples. To generate samples following $\mathbf{p^{a}_{ref}}$, we can randomly select $\mathbf{Nr}$ samples from the batch size of $\mathbf{N}$ and guide them towards male class using predictions from ${f}_{\phi}$. Similarly, for the remaining $\mathbf{N(1-r)}$ samples, we guide them towards female class. In practice, such a sample guidance follows $\mathbf{p^{a}_{ref}}$ up to some extent, but results in inferior sample quality (see Fig. 2b).
Insight. 1: Attempting to steer the generation of individual samples towards a pre-defined attribute class is overly restrictive and leads to inferior generation quality.
Remark: Sample guidance requires enforcing a preset attribute state (male / female) to each sample of a batch during the denoising process, which is too stringent and results in distorting the outputs. Given an intermediate time-stamp $\tau$, the samples in the earlier stages ($t>\tau$) of reverse diffusion process are close to the noise space, resulting in the poor classifier of the corresponding stages. The guidance becomes effective only in the later stage of denoising ($t<\tau$). However, till timestep $\tau$, some facial features are already formed for the samples [25]. Hence, enforcing a preset attribute state at later stages ($t<\tau$) is restrictive and results in sample collapse (Fig. 2-b)). For e.g., for gender attribute, if a sample has formed dominant female features till $\tau$, enforcing it to be $male$ is overly restrictive, and classifier guidance distorts the generated sample.
@@ 3.3 Distribution Guidance
Insight.2: Distribution guidance provides flexibility to batch samples to easily modify the attributes to match the defined attribute distribution.
Remark: As distribution guidance does not require presetting for attribute states for each sample, it gives more flexibility during the generation as long as $\mathbf{\hat{p}^{a}_{\theta}}$ follows $\mathbf{p^{a}_{ref}}$. E.g., in our running example of gender attribute, only those samples will change gender after $t=\tau$, which are close to the decision boundary. In contrast to sample guidance, where random samples are forced to change the attribute state (Fig.  2) resulting in inferior quality, distribution guidance is able to generate fair samples with good quality. Observe that distribution guidance covers the samples that are close to the decision boundary for gender classification, resulting in good outputs.
A major design decision for implementing distribution guidance is the selection of function $g_{\psi}$. The conventional approach is to train multiple attribute classifiers at each level of noise in the image space. Instead, we propose to leverage intermediate semantic features from the diffusion model itself and repurpose them for the task of attribute classification. This framework is super efficient as compared to image space classifiers.
@@ 3.4 Guidance in the H-space
Insight.3: H-space guidance is extremely effective and efficient as compared to image space guidance.
@@ 4.1 Evaluation Metrics
A fair generative model is evaluated on two grounds: image quality and fairness. Therefore, we begin by discussing the metrics we use to measure these two attributes of the generated images.
@@
$|\bar{p}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})|_{2}$

### v2

@@ Abstract
Diffusion Models (DMs) have emerged as powerful generative models with unprecedented image generation capability. These models are widely used for data augmentation and creative applications. However, DMs reflect the biases present in the training datasets. This is especially concerning in the context of faces, where the DM prefers one demographic subgroup vs others (eg. female vs male). In this work, we present a method for debiasing DMs without relying on additional reference data or model retraining. Specifically, we propose Distribution Guidance, which enforces the generated images to follow the prescribed attribute distribution. To realize this, we build on the key insight that the latent features of denoising UNet hold rich demographic semantics, and the same can be leveraged to guide debiased generation. We train Attribute Distribution Predictor (ADP) - a small mlp that maps the latent features to the distribution of attributes. ADP is trained with pseudo labels generated from existing attribute classifiers. The proposed Distribution Guidance with ADP enables us to do fair generation. Our method reduces bias across single/multiple attributes and outperforms the baseline by a significant margin for unconditional and text-conditional diffusion models. Further, we present a downstream task of training a fair attribute classifier by augmenting the training set with our generated data.
Code is available at - project page.
@@ 1 Introduction
Recent advancements in Diffusion Models (DM) [13, 40, 51] have garnered much interest among researchers in evaluating the quality of the generated content. These models are not only used to generate realistic content but also to augment real datasets [54, 29] for downstream tasks. However, existing DMs have been found to synthesize biased content with respect to multiple demographic factors like gender, race, etc., which can be detrimental to society once these models are deployed in the real world [35, 26, 41]. The problem is largely caused by the images used to train these models, as the outputs of these models are governed by these training datasets [35, 27]. Effects of such harmful biases have been shown by multiple recent works involving studies on DMs [35, 26, 41], GANs and other generative models [27, 6, 14]. In fact, Perera et al. [35] show that unconditional DMs–even when trained with balanced data–amplify racial biases, leading to the generation of more white-skinned faces than dark-skinned ones. Biases in the generated data are even more evident in large text-to-image DMs, e.g., models mostly tend to generate a specific gender with a given profession (like male and doctor) [28, 62, 47].
@@
Existing works on debiasing, either require a reference dataset  [6, 62] and/or allow retraining of the model [6, 60, 61]. On the contrary, our work aims to mitigate biases in both unconditional and conditional DMs, enabling fair generation without model retraining. We propose a practical setting where we provide a reference attribute distribution ($\mathbf{p^{a}_{ref}}$) for sensitive attribute $a$ and query the DM to sample images following $\mathbf{p^{a}_{ref}}$. This framework allows for adapting any existing DM to a pre-defined distribution. E.g., a user can provide $\mathbf{p^{a}_{ref}}$ as a uniform distribution for a given sensitive attribute to generate balanced attribute distribution. We believe defining $\mathbf{p^{a}_{ref}}$ provides just enough information to condition the DM for fair generation. This is an extremely practical setting for debiasing and is particularly important for large text-to-image DMs [40, 44, 38] where retraining or fine-tuning is computationally intensive.
@@
One plausible approach for fair generation is to guide every generated sample with attribute classifiers following classifier guidance [9]. However, such a framework, though simple, is overly restrictive as it requires presetting and enforcing attributes for each sample individually (which we call sample guidance). Such constraints during denoising result in inferior generation quality, as discussed in Sec.4.1. Instead, we propose to jointly denoise a whole batch of samples and guide the process with $\mathbf{p^{a}_{ref}}$ (which we call distribution guidance). Specifically, we push the generated batch attribute distribution $\mathbf{{p^{a}_{\theta}}}$ and $\mathbf{p^{a}_{ref}}$ close to each other during the reverse process. Distribution guidance provides more flexibility to each sample during generation as it does not enforce a preset of attributes on a sample basis. Intuitively, distribution guidance prioritizes transforming easier samples close to the decision boundary. This results in fair generation without sacrificing the generation quality.
@@
A major challenge for guidance-based conditioning is that it requires separate image classifiers for each noise scale of the diffusion process. To overcome this, we propose to perform guidance in a semantically rich feature space - h-space [23] of DMs. Specifically, we train an Attribute Distribution Predictor (ADP) that predicts attribute distribution directly from the h-space features. As ADP is trained on rich and discriminative h-space features, it - a) is implemented as a linear layer, b) requires minimal training data, and c) is fast in training and inference. Finally, during inference, we steer the h-space representation by matching the predictions from ADP to $\mathbf{p^{a}_{ref}}$.
@@
    A novel setting for debiasing existing DMs without retraining, given a reference attribute distribution.
@@ 2 Related Works
Biases in Generative Models. While generative models like Generative Adversarial Networks and Diffusion Models have become the de-facto tools for image generation in recent times, studies show that they are not free of biases [35, 26, 41, 27, 15, 3]. Perera et al.  [35] show that unconditional diffusion models amplify the biases in the training data with respect to gender, race, and age. Luccioni et al.  [26] identify and quantify social biases in images generated by popular text-to-image models like DALL-E 2, and Stable Diffusion v1.4 and 2. Maluleke et al.  [27] study racial biases in GANs and find that GANs mimic the racial distribution in the training data.
@@
Debiasing generative models without training. As training of GANs and DMs can be resource-consuming, many methods prefer fair generation of images without explicit training [37, 52, 29, 14]. MaGNET [14] aims to produce uniform sampling on the learned manifold of any generative model like GANs or VAEs, while Ramaswamy et al.  [37] and Tan et al.  [52] manipulate the latent space of GANs to generate balanced outputs. Controlling the generation with latent manipulation is easier in GANs due to its highly disentangled latent spaces in GANs [50, 34, 18]. GANDiffFace [29], on the other hand, generates balanced synthetic for face recognition by first generating high-quality images from different demographics using GANs and then finetuning Stable Diffusion [40] using DreamBooth [42] to generate more images of such identities with different poses, expressions, etc. Multiple works attempt to mitigate biases in vision-language models and text-conditioned diffusion models as well [7, 62, 49, 64, 4]. Chuang et al.  [7] debias the text embedding using a calibrated projection matrix and shows unbiased generations without any additional training or data. However, debiasing the unconditional DMs has received less attention, which is the main focus of this work.
@@ 3 Method
We assume a setting where we are given a pretrained DM trained on biased data and a reference distribution of the sensitive attributes $\mathbf{p^{a}_{ref}}$. Our goal is to generate data from the DM, whose generated attribute distribution $\mathbf{p^{a}_{\theta}}$ best approximates the reference $\mathbf{p^{a}_{ref}}$ without retraining. The key idea is to jointly guide the denoising of a batch of samples such that $\mathbf{p^{a}_{ref}}$approx $\mathbf{p^{a}_{\theta}}$. Directly computing $\mathbf{p^{a}_{\theta}}$ in a closed form is intractable. Instead, we train an Attribute Distribution Predictor, a linear projection that maps the intermediate batch features from h-space of a denoising network to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
@@ 3.1 Preliminary
Diffusion models have emerged as a powerful family of generative models trained to learn the data distribution by gradual denoising from a Gaussian distribution. Starting from a clean point $\mathbf{x_{0}}$, and a set of scalar values $\{\alpha_{t}\}_{t=1}^{T}$, applying $\mathbf{t}$ steps of the forward diffusion process yields a noisy data point $\mathbf{x_{t}}$, where $\bar{\alpha_{t}}$ = $\prod_{i=1}^{t}\alpha_{i}$ and
$\mathbf{x_{t}}=\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}}+(\sqrt{1-\bar{\alpha_{t}}})\epsilon,\epsilon\approx\mathcal{N}(0,I)\vspace{-2mm}$ (1)
A diffusion model is learned as a neural network $\epsilon_{\theta}$ that predicts the noise from given $\mathbf{x_{t}}$ and $\mathbf{t}$. The reverse process takes the form $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{x_{0}})$, which is parameterized as a Gaussian distribution. In this work, we consider DDIM [51] sampling which first computes an estimate of the clean data point $\mathbf{\hat{x}_{0}}$ and then sample $\mathbf{x_{t-1}}$ from $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{\hat{x}_{0}})$.
Classifier guidance is proposed to condition a diffusion model on class labels with the help of a pretrained classifier [9]. Specifically, a classifier $f_{\phi}(c|\mathbf{x_{t}},t)$ is trained on noisy images to predict the class label $c$. The gradients of the classifier are used to guide the diffusion sampling process to generate an image of the prescribed class $c$. Concretely, the classifier guidance performs sampling by updating the noise prediction $\epsilon_{\theta}(\mathbf{x_{t}},\mathbf{t})$ as follows:
@@ 3.2 Classifier guidance for debiasing
A promising approach is to leverage pretrained attribute classifiers to guide towards balanced generation. Assuming a reference attribute distribution $\mathbf{p^{a}_{ref}}$ for a binary attribute $\mathbf{a}$ (e.g., gender) and corresponding attribute classifier ${f}_{\phi}$, we parameterize $\mathbf{p^{a}_{ref}}$ as a Bernoulli distribution with parameter $\mathbf{r}$ denoting fraction of males samples. To generate samples following $\mathbf{p^{a}_{ref}}$, we can randomly select $\mathbf{Nr}$ samples from the batch size of $\mathbf{N}$ and guide them towards male class using predictions from ${f}_{\phi}$. Similarly, for the remaining $\mathbf{N(1-r)}$ samples, we guide them towards female class. In practice, such a sample guidance follows $\mathbf{p^{a}_{ref}}$ up to some extent, but results in inferior sample quality (see Fig. 3b).
Insight.1: Transforming samples close to the decision boundary is easier and results in higher quality generation.
Remark: We performed an insightful experiment for changing gender attribute (female to male) using classifier guidance. We group the samples in four quantiles based on their distance from the decision boundary of a pre-trained gender classifier in Fig. 2a). Next, we perform sample-based guidance over the samples from each quantile in Fig. 2b). The samples which are close to the decision boundary (quantiles Q1 and Q2) are easily transformed with guidance, whereas the samples away from the decision boundary (quantiles Q3 and Q4) are distorted during the guidance process. This is also quantified with the quantile-wise FID against a real set (of male images), where the samples from Q1/Q2 have better FID after conversion.
Figure 2: Distribution of samples w.r.t. decision boundary
Insight.2: Attempting to steer the generation of individual samples towards a pre-defined attribute class is overly restrictive and leads to inferior generation quality.
Remark: Sample guidance requires enforcing a preset attribute state (male/female) to each sample of a batch during the denoising process, which is too stringent and results in distorting the outputs. This is particularly bad in the case when samples from quantiles Q4 are selected for transformation, as shown in Fig. 2. Given an intermediate time-stamp $\tau$, the samples in the earlier stages ($t>\tau$) of reverse diffusion process are close to the noise space, resulting in the poor classifier of the corresponding stages. The guidance becomes effective only in the later stage of denoising ($t<\tau$). However, till timestep $\tau$, some facial features are already formed [30]. Hence, enforcing a preset attribute state at later stages ($t<\tau$) is restrictive and results in sample collapse (Fig. 3-b)). For e.g., for gender attribute, if a sample has formed dominant female features till $\tau$ it will be in Q4, and enforcing it to be $male$ is overly restrictive, resulting in distorted generations.
Figure 3: Sample guidance vs Distribution guidance. a) After a few steps of denoising ($t=\tau$), the generated samples have learned some discriminative features for gender. Sample guidance randomly selects samples from the batch uniformly from all the quantiles for conversion enforcing samples with dominant female features to also change. However, distribution guidance majorly converts the samples close to the decision boundary (Q1/Q2), which is easy to convert. b) Visualization of the generated samples. Samples transformed with distribution guidance are natural looking without any distortion, whereas images with sample guidance suffer from distortion or unnatural appearance.
@@ 3.3 Distribution Guidance
Insight.3: Distribution guidance provides flexibility to batch samples and transforms easier samples close to the decision boundary to match the required distribution.
Remark: As distribution guidance does not require presetting for attribute states for each sample, it gives more flexibility during the generation as long as $\mathbf{\hat{p}^{a}_{\theta}}$ follows $\mathbf{p^{a}_{ref}}$. E.g., in our running example of gender attribute, only those samples will change gender after $t=\tau$, which are close to the decision boundary (in Q1 & Q2). In contrast, in sample guidance, random samples are forced to change the attribute state (Fig. 3) and are spread equally in all four quantiles, resulting in inferior quality. This is quantified by the fraction of samples being transformed ($f_{Qi}$) with each guidance in Fig. 3a), where distribution guidance majorly transforms samples from Q1 & Q2 and results in fair and high-quality samples (Fig. 3b)).
A major design decision for implementing distribution guidance is the selection of function $g_{\psi}$. The conventional approach is to train multiple attribute classifiers at each level of noise in the image space. Instead, we propose to leverage intermediate semantic features from the diffusion model itself and repurpose them for the task of attribute classification. This framework is very efficient as compared to image space classifiers.
@@ 3.4 Guidance in the H-space
Insight.4: H-space guidance is extremely effective and efficient as compared to image space guidance.
@@ 4.1 Evaluation Metrics
A fair generative model is evaluated on two grounds: image quality and fairness. We discuss the metrics used to measure these two aspects of the generated images below.
@@
$||\bar{p}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})||_{2}$